# PIA: Your Personalized Image Animator via Plug-and-Play Modules in Text-to-Image Models

### 21 Dec 2023 v1

@@ 4.3 Analysis
The condition module enables PIA to borrow appearance features from the conditional frame so that temporal alignment layers can focus more on motion generation. These designs enable PIA to perform better image and text alignment in generated videos. The strong prompt control ability of PIA can be seen in Fig. 6. During the test, we observed that a simple prompt design produces great visual effects, as shown in the figure. PIA is able to generate fancy results by adding new elements to the generated videos with smooth motion quality based on text prompts. Unlike Pika Labs or Gen-2 prone to generate small motion within 3 seconds, PIA generates video with more text alignment motion.
@@
PIA can control the magnitude of motion by adjusting the frame affinity. In the inference stage, we constructed three types of frame affinity patterns based on real video data, and the test results are shown in Fig. 7. Specifically, we use the statistics in the training stage, to construct the similarity affinity pattern. We broadly divide the range of motion into three categories and provide the corresponding affinity in inference time. In Fig. 7, the frame affinity gradually decreases from top to bottom, indicating an increasing magnitude of motion. From the results, it can be observed that the generated motion aligns with our expectations, demonstrating the effectiveness of the similarity channel in controlling motion magnitude. Meanwhile, PIA can also control the position of the input image in the video by adjusting the Similarity channel. In short, by using different frame affinity during the inference stage, users can flexibly control the motion magnitude and conditional frame position in the generated video, leading to a more controllable animation.
@@
So far, we have discussed using models and input images in the same domain to generate videos. However, we have found interesting facts that if the two have different domains, such as providing a real-style image and using an anime-style model to generate motion, PIA can achieve style transfer effects. As shown in Fig. 8. Frame affinity helps generate videos with smooth style transfer between image style and model style. This feature of PIA allows users to have a better-personalized experience.
Figure 8: Style transfer. PIA is able to achieve style transfer in generated videos by applying a personalized text-to-image model with a different domain style to the given image.
Effectiveness of Frame Affinity.
In the training stage, we calculate the similarity score for affinity hints as mentioned before. We perform an ablation study with one-hot frame affinity to prove the effectiveness of similarity frame affinity. One-hot frame affinity is constructed by 1 in the condition frame and 0 in the other frames which is able to indicate the position of the condition frame. We evaluate two methods on the AnimateBench. One-hot frame affinity obtains the image CLIP score of 210.9, while the similarity frame affinity reaches 225.9. Moreover, we also show the visual result of video frames generated by the two constructions in Fig. 9. We can find that frame affinity constructed by similarity allows the model to learn identity better. Note that in different constructions, we use the same pattern of affinity in training and inference.
Figure 9: Effectiveness of frame affinity. With frame affinity, PIA is able to generate videos with stronger image alignment.
@@
PIA is capable of animating personalized images with high text and image alignment with text conditions. However, we have observed that videos generated by PIA exhibit color discrepancy when applied to images with significantly different styles from the training data, as shown in Fig. 10. We hypothesize the color discrepancy is caused by the WebVid [2] domain. PIA learns the color distribution of WebVid, and when there is a notable difference between the personalized image during inference, the color discrepancy may arise. Additionally, when generating images, specific trigger words such as a certain style may be required. However, when generating videos, if the prompts do not include these words, it can also result in color discrepancy. We further speculate that training on a more diverse range of video data with various styles and content and providing complete trigger words during inference could mitigate this phenomenon.
@@ 5 Conclusions
In this paper, we introduce PIA  a powerful solution for personalized image animation. Our method shows excellent image alignment and motion controllability, ensuring text-based customized animation for users. To achieve high alignment, we introduce condition module and inter-frame affinity. They endow PIA borrow appearance features from the conditional frame, leading temporal layers to focus more on motion generation. Furthermore, we construct an animation benchmark named AnimateBench to evaluate the animation performance of PIA and other methods. Experimental results demonstrate that PIA performs excellent on image animation tasks and shows various interesting extended applications.
In future work, we first plan to extend our method to a more powerful base text-to-image model, e.g., SDXL [26], which is becoming the new trend in the Stable Diffusion community. Second, we plan to study the limitation of color shift caused by the low-quality training dataset.

### 22 Mar 2024 v2

@@ 4.3 Analysis
The condition module enables PIA to borrow appearance from the condition frame so that temporal alignment layers can focus more on motion generation. These designs enable PIA to perform better image and text alignment in generated videos. The strong prompt control ability of PIA can be found in Fig. 6. During the test, we observed that a simple prompt design produces great visual effects, such as fire and lightning. PIA can generate fancy results by adding new elements to the generated videos with smooth motion quality based on text prompts. Unlike Pika Labs or Gen-2 which are prone to generate small motion while ignoring text prompt, PIA generates video with more text alignment motion.
@@
PIA can control the magnitude of motion by adjusting the frame affinity. Specifically, we first make statistics of the frame affinity across the whole video training dataset. Then, according to the statistical results, we roughly divide the range of frame affinity into three groups and find three corresponding sets of inter-frame affinity values. Finally, during inference, given the same condition image and text prompt, PIA can achieve videos with different motion magnitudes by using different sets of inter-frame affinity inputs. As shown in Fig. 7, the frame affinity gradually decreases from top to bottom, indicating an increasing magnitude of motion. From the results, it can be observed that the dog can jump with different magnitudes, demonstrating the effectiveness of the frame affinity in controlling motion magnitude. In short, by using different frame affinity during the inference stage, users can flexibly control the motion magnitude of the generated video, leading to a more controllable animation.
@@
So far, we have discussed using models and input images in the same domain to generate videos. However, we have found interesting facts that if the two have different domains, such as providing a real-style image and using an anime-style model to generate motion, PIA can achieve style transfer effects. As shown in Fig. 8, frame affinity helps generate videos with smooth style transfer between image style and model style. This feature of PIA allows users to have a better personalized experience.
Figure 8: Style transfer. Given the same input image and text prompt, PIA can generate different styles of videos like style transfer by using different personalized text-to-image models (e.g., RCNZ 3d Cartoon and ToonYou in this case).
Ablation Study on Frame Affinity.
During training, we calculate the similarity between the condition frame and each frame as the affinity hints. To verify the effectiveness of the inter-frame affinity design, we conduct an ablation study by comparing it with one-hot frame affinity. Specifically, one-hot frame affinity is constructed by setting 1 in the condition frame and setting 0 in the other frames. Such a design can indicate the position of the condition frame. We evaluate two methods on the AnimateBench. One-hot frame affinity obtains the image CLIP score of 210.9, while the inter-frame affinity reaches 225.9. Moreover, we show the visual result of video frames generated by these two designs in Fig. 9. We can find that frame affinity constructed by similarity allows the model to align with the condition frame better. Note that in different constructions, we use the same pattern of affinity in training and inference.
Figure 9: Ablation study on the design of inter-frame affinity. Compared with the one-hot condition design, the design of inter-frame affinity in PIA shows better image alignment.
@@
First, we have observed that videos generated by PIA exhibit color discrepancy when applied to images with significantly different styles from the training data, as shown in Fig. 10. We hypothesize the color discrepancy is caused by training on the WebVid dataset [2]. WebVid consists mainly of real-world recordings, characterized by frames with motion blur and compression artifacts [2, 12]. Hence, the quality and diversity of these frames are lower than image datasets that include professional photography and artistic paintings [4, 12]. Training with this noticeable quality domain gap can inevitably limit the animation of complex images with significant stylistic differences.
Secondly, in personalized image generation, users rely on specific trigger words in text prompts to evoke desired styles. However, during image animation, if these trigger words are absent from the animation prompts, significant color inconsistencies may arise. We further speculate that training the model on a wider range of video data encompassing diverse styles and content can ease this issue. Additionally, ensuring the inclusion of complete trigger words during inference could help mitigate this phenomenon.
@@ 5 Conclusions
In this paper, we introduce PIA  a powerful solution for personalized image animation. Our method shows excellent image alignment and motion controllability, ensuring text-based customized animation for users. Furthermore, we construct an animation benchmark named AnimateBench to evaluate the animation performance of PIA and other methods. Experimental results demonstrate that PIA performs excellent on image animation tasks and shows various interesting extended applications.
6 Acknowledgement
This project is supported by the National Key R&D Program of China (No. 2022ZD0161600)